7) A number decreased by 2 is 40% of that number increased by 7.what is the number?
Answer #1

Let X be the unkown number. From the given information, we can formulate the linear equation as,

(X-2)=\frac{40}{100}*(X+7)

Hence, on simplification,

100X - 200 = 40X + 280

100X - 40X = 200 + 280

60X = 480

Hence,

X = 480/60

X = 8

Hence the unknown number is 8.